A ball is projected vertically upward with a velocity of 80m/s. Two seconds later a second ball is projected vertically upward with a velocity of 60m/s. At what point above the surface of earth will they meet?


1
Expert's answer
2021-03-10T17:21:30-0500

Let's take the upwards as the positive direction. After "t=2\ s" the first ball reached the height "y(2)" and its vertical displacement can be written as follows:


"y_1=y(2)+v(2)t-\dfrac{1}{2}gt^2."


Let's find the velocity and height of the first ball after "t=2\ s":


"v=v_0-gt=80\ \dfrac{m}{s}-9.8\ \dfrac{m}{s^2}\cdot2\ s=60.4\ \dfrac{m}{s},""y(t=2\ s)=80\ \dfrac{m}{s}\cdot2\ s-\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(2\ s)^2=140.4\ m."


Then, we can write the vertical displacement of the second ball projected vertically upward two seconds later:


"y_2=v_{02}t-\dfrac{1}{2}gt^2"


At the moment when they meet "y_1=y_2" and we can write:


"y(2)+v(2)t-\dfrac{1}{2}gt^2=v_{02}t-\dfrac{1}{2}gt^2."


Substituting "y(2)" and "v(2)" in the previous equation, we get:


"140.4+60.4t-\dfrac{1}{2}gt^2=60t-\dfrac{1}{2}gt^2,""140.4=-0.4t""t=-351\ s."

The time is negative, which means that two balls never meet at the height above the earth.